Determine the molatility of a solution made by dissolving 52.6 g of LiF in 856 g of H2O.
Answer #1

molality is no. of moles of solute in 1 kg of solvent

here moles of solute LiF = mass / molar mass = 52.6 / 26 = 2.02 mol

mass of solvent = 856 g

hence 856 g of water contain 2.02 mol of LiF

therefore 1 kg or 1000 g of water will have (2.02 / 856) * 1000 = 2.36 mol

hence molality = 2.36 m

PLEASE LEAVE THE POSITIVE FEEDBACK IF YOU LIKE THE SOLUTION